create new tag
view all tags

Laser-wire Alignment

Laser-wire system was aligned in preparation for shift this week. Laser beam path was aligned over weekend using first alignment laser then pulsed laser. The following measurements were used to align the laser beam to be perpendicular to the chamber as well as centred on the aperture.

  • Back reflection from chamber + iris
  • Back reflection from mirror in lens mount + iris

This also ensured that the lens mount was parallel to the vacuum window.

OTR Alignment

The OTR alignment laser was aligned along the electron beam pipe and observed using the OTR screen at the LWIP. The OTR screen was moved vertically to still show the reflected alignment laser (red) but allow some of the pulsed laser light underneath (the LW lens was removed for this). The angle of the screen was adjusted to overlap these at a distance of approximately 1m.

  • Manipulator Angle: 0.275 degrees

The OTR alignment laser was then used to construct the OTR beam path to the camera. The polariser rotation mount only just fits in the newly rotate OTR housing on the side of the LW housing. However, the cable protruding from it does not - will either have to make a hole in the enclosure or move the mounts so that the support rail for the OTR is on the short side of the housing. This would require a full realignment of the OTR beam line. The camera and filter wheel are installed and connected. The camera has been tested, but the filter wheel has not.

The OTR switching stage was tested successfully and its positions are

  • LW main line - 0
  • OTR line - 100000

APD Dump

The APD dump stage was also tested. This stage (Thorlabs MTS50) does not home. The limit switches work when pressed - the software recognises this, but these are not triggered when the stage passes over them causing the stage to grind into the ends. The stage was manually backed off from the end by a couple of milimetres and then power cycled to zero the encoder.

  • Do not attempt to home the APD stage

The positions for operation are

  • APD open - 0
  • APD closed - 43

APD Testing & Alignment

The APD was connected to one of the heliax cables and switched on. The OTR switch stage was set to 0 for the main LW line. Laser was set to RGA only and with lots of attenuation - only barely visible to the naked eye. The APD dump was opened. Using the network Zscope we now have, the APD signal was observed. Normal operation of the APD:

  • Around 200mV signal with RGA only + full rack of attenuators

No signal was observed intially. A copy of the gate signal for the Cherenkov signal was used (no spare gate generators in DAQ) as this would be roughly the right time. It was however approximately 100ns off - too far away from the APD signal to observe it on a sensible timebase for the ~1ns long APD signal. When a signal was observed, the scope was triggered from it itself. The alignment of the APD was optimised in the following order:

  • XY position (micrometers)
  • XY Angle (screws on mount)

  • ~200mV signal was observed with all attenuators available in laser beam without saturation.

The stage for the APD dump was manouvered so that the ceramic shield covering the APD moved parallel to the front surface of the collecting cylinder but didn't scrape against it. It was positioned to be as close as possible.

Laser Referencing

The OTR screen was lowered into the chamber until the laser beam was interrupted. The laser was observed on a card after the recollimating lens post-IP.

  • Laser Position between -35650 and -35670

In between these positions the laser beam goes from fully observable to fully occluded. Although a 20 micron range, the screen is ~300 microns thick so will cover a large part of the laser waist. The screen position could be moved to minimise this, but this would align the back side of the screen facing away from the laser beam with the focus and not the side that will hit the electron beam. With ~20 microns distance to occlude the beam, the screen must be very close to the focus of the laser beam though.

BPM Connection

To provide a trigger very close to the beam arrival time at the LWIP, the button BPM attached to our chamber upstream of the interaction point was connected. The 8 BPM clables in our main cable bundle don't quite reach to the new LWIP. Therefore, a 2m extension was borrowed from YoungIm Kim to extend one of them. It must be returned next weekend though. It was connected to the bottom vertical BPM output.

Laser Energy & Dumping

No stable solution exists yet for measuring the laser energy each pulse. One of the high damage threshold energy meters along with the attenuator attached was placed in the LW main line as a beam dump. It is currently not connected to anything. It was the only suitable beam dump available though.

It is intended to find a solution to laser energy measurements during the week and install a permanent solution next weekend. Possible solutions:

  • Try S-Link2 triggered energy meter over USB - readout problems were over network
  • If this works - put low energy head behind first turning mirror and calibrate attenuation
  • Photodiode behind mirror - calibrate / check linearity - use ADC to digitise

Edit | Attach | Watch | Print version | History: r3 < r2 < r1 | Backlinks | Raw View | Raw edit | More topic actions

Physics WebpagesRHUL WebpagesCampus Connect • Royal Holloway, University of London, Egham, Surrey TW20 0EX; Tel/Fax +44 (0)1784 434455/437520

Topic revision: r3 - 05 Dec 2011 - LauraCorner

This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2023 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding RHUL Physics Department TWiki? Send feedback